# Neural Networks Learn Statistics of Increasing Complexity

### v2

@@ 1 Introduction
Figure 1: (left) Pekinese dog image from the ImageNet training set. (center) Image after quantile normalizing its pixels to match the marginal distribution of the goldfish class on ImageNet. The grass is now a slightly darker shade of green and the dog’s fur has a reddish hue. (right) Synthetic “goldfish” generated by sampling pixel independently from its marginal distribution on ImageNet.
@@ 2 Theory and Methods
where $\alpha$ is a multi-index, or a $d$-tuple assigning an integer to each coordinate of $\mathbf{x}$. Recall that taking a vector to the power of a multi-index denotes a product of the components of the vector, where each component of the index indicates the multiplicity: e.g. if $\alpha=(1,4,6)$, the expression $(\mathbf{x}-\boldsymbol{\mu})^{\alpha}$ denotes the product $(x_{1}-\mu_{1})(x_{2}-\mu_{2})^{4}(x_{3}-\mu_{3})^{6}$. Similarly, ($\partial^{\alpha}\mathcal{L}$) is shorthand for the mixed partial derivative $\frac{\partial^{11}}{\partial x_{1}\partial x_{2}^{4}\partial x_{3}^{6}}\mathcal{L}$. The factorial $\alpha!$ denotes the product of the factorials of the components: 1!×4!×6!=172801$\times$ 4$\times$ 6!=172801 ! × 4 ! × 6 ! = 17280.
@@ 2.3 Optimal Transport
We operationalize Criterion 1 using optimal transport (OT) theory, which provides tools for transforming samples from one probability distribution into samples from another while minimizing the average distance that samples are moved. We use two OT methods in our experiments: coordinatewise quantile normalization, which primarily affects the first order moments of the distribution, and Gaussian OT, which affects both the first and second-order moments.
@@
Given two Gaussians $P=\mathcal{N}(\boldsymbol{\mu}_{P},\boldsymbol{\Sigma}_{P})$ and $Q=\mathcal{N}(\boldsymbol{\mu}_{Q},\boldsymbol{\Sigma}_{Q})$ supported on $\mathbb{R}^{d}$, the map $T(\mathbf{x})=\mathbf{A}(\mathbf{x}-\mathbf{m}_{P})+\mathbf{m}_{Q}$ is the optimal transport map from $P$ to $Q$ under the squared Euclidean cost function, where
@@ 2.4 Maximum Entropy Sampling
Famously, the maximum entropy distribution supported on $\mathbb{R}^{d}$ with known mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$ is the Gaussian distribution $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$. We therefore use Gaussians in many of our experiments in Sec 3. In addition to Gaussians, we use another second-order method (hypercube constrained sampling) and a first-order method (independent coordinate sampling), both of which we describe below.
@@
One problem with using Gaussians to generate synthetic images is that natural images are constrained to a hypercube: RGB pixel intensities are in the range [0,255][0,255][ 0 , 255 ], but nonsingular Gaussian distributions assign positive probability density to all of $\mathbb{R}^{d}$, so that a typical sample will often lie outside the hypercube of natural images. [^9] Strictly speaking, Gaussians also violate the assumption of compact support that we made earlier. In high dimension, though, almost all the probability mass of a Gaussian is contained in the typical set, a compact region near the boundary of an ellipsoid surrounding the mean (Carpenter, 2017). We might expect neural networks to quickly adapt to such a simple box constraint on the support, so we would ideally like to subject our synthetic images to this constraint.
In the 1D case, the maximum entropy probability density with known mean and variance supported on a finite interval $[a,b]$ has the form $p(x)=\exp(-\lambda_{0}-\lambda_{1}x-\lambda_{2}x^{2})$ (Dowson & Wragg, 1973), but we are unaware of any such analytical formula for the multidimensional case. [^10] For some values of the Lagrange multipliers, the Dowson & Wragg (1973) formula corresponds to a truncated normal distribution. For sufficiently large variances, however, the density takes on a U-shape. Natural image pixels often have bimodal distributions with high variance, and the multidimensional analogue of the U-density may be complex and difficult to sample from directly.
@@
For many sets of constraints, there is no known closed-form solution for the density, precluding standard sampling techniques like Markov chain Monte Carlo. For these cases, we propose a novel technique for approximate sampling: use gradient-based optimization to directly produce a finite set of samples whose statistics match the desired ones, while maximizing the Kozachenko-Leonenko estimate for the entropy of the implicit population distribution (Kozachenko & Leonenko, 1987; Sablayrolles et al., 2018). See Appendix E for implementation details and Appendix F for a discussion of computational requirements.
@@
By constraining the coordinatewise marginals, we ensure that all of the homogeneous moments, or moments of the form $\mathbb{E}[(x_{i})^{n}]$, match those of the true data distribution, while the mixed moments, e.g. $\mathbb{E}[x_{i}x_{j}]$ for $i\neq j$, will generally not match. In high dimension, almost all moments of order greater than one are mixed rather than homogeneous, so ICS matches the first order moments and almost none of the higher order ones. This is useful because, to our knowledge, there is no analytical formula for the maximum entropy distribution given only a mean and box constraint.
@@ 2.5 Discrete Domains
[…]
@@ 3 Image Classification
Because Gaussian optimal transport (Sec. 2.3) requires $O(d^{3})$ compute and $O(d^{2})$ memory, [^12] The covariance matrix has $d^{2}$ elements, where $d$ is the number of pixels, so it is actually $O(n^{4})$ in the height or width dimension of the image. We also ran into a software limitation in early experiments where NumPy and PyTorch eigensolvers would crash when fed the very large covariance matrices produced by high-resolution image datasets; see PyTorch issue #92141 for discussion. we focus on datasets with $\times 32$ 32 or $\times 64$ 64 resolution images for our primary experiments. Specifically, we examine the popular image classification datasets CIFAR-10 (Krizhevsky et al., 2009), Fashion MNIST (Xiao et al., 2017), MNIST (LeCun et al., 1998), and SVHN (Netzer et al., 2011).
We also build a new image classification dataset, CIFARNet, consisting of 200K images at $\times 64$ 64 resolution sampled from ImageNet-21K, using ten coarse-grained classes that roughly match those of CIFAR-10. The larger number of images per class (20K) allows us to get a good estimate of the class-conditional covariance matrices needed for Gaussian optimal transport, which at this resolution contain $(3\times 64\times 64)^{2}\approx 1.5\times 10^{8}$ entries each. See Appendix B for more details on CIFARNet.
@@
Figure 6: Cross entropy loss of Pythia suite evaluated on unigram and bigram sequences (N=1024). For comparison, the Shannon entropy of the unigram distribution is 2.89 bits per byte (bpb), and is 2.04 bpb for the bigram distribution. Both plots exhibit “double descent” scaling: loss reaches a trough between $2^{6}$ and $2^{8}$ steps, increases until $2^{10}$ steps, then decreases again as the model learns to match the data generating process in-context (Fig. 7).
@@ 3.3 Results
We include three different conditions in our maximum entropy sampling experiments: 1st order (ICS), 2nd order (Gaussian sampling), and 2nd order plus a hypercube constraint.
@@
Overall, we found that model scale has a remarkably small effect on the learning curves, so we display curves averaged over scales in bold, with individual model scales shown as translucent lines. Each curve is itself averaged over three random seeds.
@@ 4 Language Modeling
Figure 7: In-context learning of the maximum entropy bigram sequence occurs after step 1,000 in Pythia 12B. Some in-context learning of the maximum entropy unigram sequence occurs by step 1000, with more at step 143,000.
@@
While we include skip-grams (e.g. the     dog) in our formal definition of $n$-gram frequency (Def. D.1), we do not include them in these experiments for tractability reasons: they would greatly increase the memory and storage requirements of maximum entropy sampling. We hope to explore the effect of skip-gram statistics in future work.
@@
We compute token unigram and bigram frequencies across Pythia’s training corpus, the Pile (Gao et al., 2020), and use these statistics to construct maximum entropy $n$-gram language models. We autoregressively sample sequences of length 2049 from the $n$-gram LMs, and evaluate Pythia’s cross entropy loss on these maximum entropy samples at each checkpoint.
Additionally, we evaluate Pythia 12B’s cross entropy loss over each token position in the maximum entropy $n$-gram sequences for four different checkpoints in order to detect the development of in-context learning. If in-context learning is involved in making predictions for these sequences at a training step, cross entropy loss should decrease over successive token positions in the sequence for that step (Olsson et al., 2022).
@@ 4.1 Results
Across all models, we observe non-monotonicity in the $n$-gram sequence loss curves, where loss steeply reduces and then increases to a lesser extent. However, unlike in the image classification tasks, the loss reverts to a monotonic regime later in training. We hypothesize that this is caused by the development of in-context learning sufficient to improve n-gram sequence predictions. We observe correlational evidence in the $n$-gram sequence loss over increasing token indices and training steps in Pythia 12B (Fig. 7), where in-context learning seems to emerge in the same training step where the non-monotonic regime ends. Fascinatingly, smaller models seem to resume the standard ‘U’-shaped loss pattern in the later portions of training. [^15] Arguably, this pattern applies to models of all sizes on unigram sequences, but the tiny increases in loss for the larger models are within the margin of error for these experiments.
@@ 5 Conclusion
We propose two criteria that operationalize what it means for models to exploit moments of a given order, then describe methods of generating synthetic data that test whether a network satisfies both criteria, using theoretically grounded approaches relying on optimal transport theory and the principle of maximum entropy. We extend our analysis to discrete sequences by proving a novel equivalence between $n$-gram statistics and statistical moments.
We find new compelling empirical evidence that neural networks learn to exploit the moments of their input distributions in increasing order, and further find “double descent” in the degree to which language model predictions match low-order data statistics, driven by in-context learning on longer sequences. Our contributions strengthen the case for the distributional simplicity bias (DSB), refine our understanding of how DSB influences early learning dynamics, and provide a foundation for further investigations into DSB.
5.1 Acknowledgements
@@ Appendix B CIFARNet dataset
We ensured class balance by randomly sampling 20K images from each synset. Images were directly resized to $\times 64$ 64 resolution without center cropping.
@@ C.5 Street View Housing Numbers
Figure 11: The Street View Housing Numbers dataset is somewhat of an outlier in that none of the models ever exceed random baseline accuracy on 1st order synthetic images. We hypothesize this is because of the extreme diversity of colors, fonts, and background textures in SVHN, which make “simple” first order features less discriminative for classifying digits. We also found it necessary to use a smaller learning rate to achieve convergence on this dataset (Footnote 13).
@@ C.6 Pythia Language Models
Figure 12: (top) Average cross entropy loss of Pythia models evaluated on unigram and bigram sequences, (bottom) KL divergence between the predictions of our $n$-gram language models and the predictions of Pythia checkpoints when evaluated on chunks of tokens sampled from the Pile ($N=1024$.)
@@
Figure 13: Effects of fast and slow learning rate warmup on n-gram sequence loss and KL divergence, Pythia 14M and 70M.
@@
Figure 14: N-gram sequence loss and KL divergence over Pythia sizes and random seeds (1024 sequences sampled at each step, 9 seeds for Pythia 14M, 70M and 160M, 4 seeds for Pythia 410M.)
Appendix D Discrete domain proofs

### v3

@@ 1 Introduction
Figure 1: (left) Pekinese dog image from the ImageNet training set. (center) Image after quantile normalizing its pixels to match the marginal distribution of the goldfish class on ImageNet. The grass is now a slightly darker shade of green and the dog’s fur has a reddish hue. (right) Synthetic “goldfish” generated by sampling each pixel independently from its marginal distribution.
@@ 2 Theory and Methods
where $\alpha$ is a multi-index, or a $d$-tuple assigning an integer to each coordinate of $\mathbf{x}$. Recall that taking a vector to the power of a multi-index denotes a product of the components of the vector, where each component of the index indicates the multiplicity: e.g. if $\alpha=(1,4,6)$, the expression $(\mathbf{x}-\boldsymbol{\mu})^{\alpha}$ denotes the product $(x_{1}-\mu_{1})(x_{2}-\mu_{2})^{4}(x_{3}-\mu_{3})^{6}$. Similarly, ($\partial^{\alpha}\mathcal{L}$) is shorthand for the mixed partial derivative $\frac{\partial^{11}}{\partial x_{1}\partial x_{2}^{4}\partial x_{3}^{6}}\mathcal{L}$. The factorial $\alpha!$ denotes the product of the factorials of the components: $1!\times 4!\times 6!=17280$.
@@ 2.3 Optimal Transport
We operationalize Criterion 1 using optimal transport (OT) theory, which provides tools for transforming samples from one probability distribution into samples from another while minimizing the average distance that samples are moved. We use three OT methods in our experiments: coordinatewise quantile normalization and bounded shift, which primarily affect the first order moments of the distribution, and Gaussian OT, which affects both the first and second-order moments.
@@
Given two Gaussians $P=\mathcal{N}(\boldsymbol{\mu}_{P},\boldsymbol{\Sigma}_{P})$ and $Q=\mathcal{N}(\boldsymbol{\mu}_{Q},\boldsymbol{\Sigma}_{Q})$ supported on $\mathbb{R}^{d}$, the map $T(\mathbf{x})=\mathbf{A}(\mathbf{x}-\mathbf{m}_{P})+\mathbf{m}_{Q}$ is the optimal transport map from $P$ to $Q$ under the L2 cost function, where
@@
Bounded Shift
In Appendix D, we derive an algorithm for shifting the mean of an empirical distribution to a desired value, keeping its support constrained to the interval $[0,1]$, and minimizing the transport cost. We use this algorithm to graft the mean of one class onto another, while ensuring the pixel intensities of the edited images are valid.
@@ 2.4 Maximum Entropy Sampling
Famously, the maximum entropy distribution supported on $\mathbb{R}^{d}$ with known mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$ is the Gaussian distribution $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$. We therefore use Gaussians in many of our experiments in Sec 3. In addition to Gaussians, we use hypercube constrained sampling to generate synthetic samples using second and third order statistics, and we use two first-order methods (Conrad sampling and independent coordinate sampling). We explain these methods below.
@@
One problem with using Gaussians to generate synthetic images is that natural images are constrained to a hypercube: RGB pixel intensities are in the range $[0,255]$, but nonsingular Gaussian distributions assign positive probability density to all of $\mathbb{R}^{d}$, so that a typical sample will often lie outside the hypercube of natural images. [^9] Strictly speaking, Gaussians also violate the assumption of compact support that we made earlier. In high dimension, though, almost all the probability mass of a Gaussian is contained in the typical set, a compact region near the boundary of an ellipsoid surrounding the mean (Carpenter, 2017). We might expect neural networks to quickly adapt to such a simple box constraint on the support, so we would like to subject our synthetic images to this constraint.
In the 1D case, the maximum entropy probability density with known mean and variance supported on a finite interval $[a,b]$ has the form $p(x)=\exp(-\lambda_{0}-\lambda_{1}x-\lambda_{2}x^{2})$ (Dowson & Wragg, 1973), [^10] For some values of the Lagrange multipliers, the formula corresponds to a truncated normal distribution. For sufficiently large variances, the density takes on a U-shape. but we are unaware of an analytic formula for the Lagrange multipliers in the multidimensional case, making it intractable to solve. [^11] The log-density of the multidimensional max entropy distribution must be a quadratic form, just like the multivariate normal, but the “scale” matrix may not be p.s.d., and solving for the parameters seems challenging.
We prove in Theorem E.1 that the maximum entropy distribution supported on $[0,1]$ with a fixed mean $\mu\neq\frac{1}{2}$ and unconstrained variance takes the form $p(x)=\frac{b\exp(-bx+b)}{\exp(b)-1}$, where the parameter $b$ can be found using Newton’s method. This formula is not well-known, although an alternative derivation can be found in Conrad (2004). To isolate the effect of first-order statistics, we first fit a Conrad distribution to the mean of each coordinate of the images. We then generate synthetic images using inverse transform sampling to produce a value for each coordinate independently.
@@
For many sets of constraints, there is no known closed-form solution for the density, precluding sampling techniques like Markov chain Monte Carlo. For these cases, we propose a novel technique for approximate sampling: use gradient-based optimization to directly produce a finite set of samples whose statistics match the desired ones, while maximizing the Kozachenko-Leonenko estimate for the entropy of the implicit population distribution (Kozachenko & Leonenko, 1987; Sablayrolles et al., 2018). See truncated_normal.py in our codebase for implementation details and Appendix G for a discussion of computational and memory requirements.
@@
By constraining the coordinatewise marginals, we ensure that all of the homogeneous moments, or moments of the form $\mathbb{E}[(x_{i})^{n}]$, match those of the true data distribution, while the mixed moments, e.g. $\mathbb{E}[x_{i}x_{j}]$ for $i\neq j$, will generally not match. In high dimension, almost all moments of order greater than one are mixed rather than homogeneous, so ICS matches the first order moments and almost none of the higher order ones.
@@ 2.5 Discrete Domains
[$n$-gram statistics are moments] Let $\mathcal{V}^{N}$ be the set of token sequences of length $N$ drawn from a finite vocabulary $\mathcal{V}$, let $P$ be a distribution on $\mathcal{V}^{N}$, and let $f:\mathcal{V}^{N}\rightarrow\{0,1\}^{N\cdot|\mathcal{V}|}$ be the function that encodes a length-$N$ sequence of tokens as a flattened concatenation of $N$ one-hot vectors of dimension $|\mathcal{V}|$. Let $f_{\sharp}P$ be the pushforward of $P$ through this one-hot encoding, i.e. its analogue in $\{0,1\}^{N\cdot|\mathcal{V}|}$.
@@ 3 Image Classification
Because Gaussian optimal transport (Sec. 2.3) requires $O(d^{3})$ compute and $O(d^{2})$ memory, [^13] The covariance matrix has $d^{2}$ elements, where $d$ is the number of pixels, so it is actually $O(n^{4})$ in the height or width dimension of the image. We also ran into a software limitation in early experiments where NumPy and PyTorch eigensolvers would crash when fed the very large covariance matrices produced by high-resolution image datasets; see PyTorch issue #92141 for discussion. we focus on datasets with $32\times 32$ or $64\times 64$ resolution images for our primary experiments. Specifically, we examine the popular image classification datasets CIFAR-10 (Krizhevsky et al., 2009), Fashion MNIST (Xiao et al., 2017), MNIST (LeCun et al., 1998), and SVHN (Netzer et al., 2011).
We also build a new image classification dataset, CIFARNet, consisting of 200K images at $64\times 64$ resolution sampled from ImageNet-21K, using ten coarse-grained classes that roughly match those of CIFAR-10. The larger number of images per class (20K) allows us to get a good estimate of the class-conditional covariance matrices needed for Gaussian optimal transport, which at this resolution contain $(3\times 64\times 64)^{2}\approx 1.5\times 10^{8}$ entries each. See Appendix B for more details on CIFARNet.
@@
Figure 6: Cross-entropy loss of Pythia suite evaluated on 1- through 4-gram sequences ($N=4096$.) For comparison, the Shannon entropy of the 1-gram distribution is 2.89 bits per byte (bpb), and is 2.04 bpb for the 2-gram distribution. The first two plots exhibit “double descent” scaling: loss reaches a trough between $2^{6}$ and $2^{8}$ steps, increases until $2^{10}$ steps, then decreases again as the model learns to match the data generating process in-context (Fig. 7).
@@ 3.3 Results
We include four different conditions in our maximum entropy sampling experiments: 1st order (ICS), 2nd order (Gaussian sampling), and both 2nd and 3rd order plus a hypercube constraint.
@@
Overall, we found that model scale has a remarkably small effect on the learning curves, so we display curves averaged over scales in bold, with individual model scales shown as translucent lines.
@@ 4 Language Modeling
Figure 7: Mean loss over n-gram sequence positions at six Pythia-12B training steps. In-context learning of the maximum entropy bigram sequence occurs after step 8,000. Some in-context learning of the maximum entropy unigram sequence occurs by step 1,000, with more after step 8,000.
@@
While we include skip-grams (e.g. the       dog) in our formal definition of $n$-gram frequency (Def. F.1), we do not include them in these experiments for tractability reasons: they would greatly increase the memory and storage requirements of maximum entropy sampling. We hope to explore the effect of skip-gram statistics in future work.
@@
We compute token unigram and bigram frequencies across Pythia’s training corpus, the Pile (Gao et al., 2020), and use these statistics to construct maximum entropy $n$-gram language models. We autoregressively sample sequences of length 2049 from the $n$-gram LMs, and evaluate Pythia’s cross-entropy loss on these maximum entropy samples at each checkpoint. We repeat the procedure for 3- and 4-grams using a subset of Pythia’s training corpus.
Additionally, we evaluate Pythia 12B’s cross-entropy loss over each token position in the maximum entropy $n$-gram sequences for six different checkpoints in order to detect the development of in-context learning. If in-context learning is involved in making predictions for these sequences at a training step, cross-entropy loss should decrease over successive token positions in the sequence for that step (Olsson et al., 2022).
@@ 4.1 Results
Across all models, we observe non-monotonicity in the unigram and bigram sequence loss curves, where loss steeply reduces and then increases to a lesser extent. However, unlike in the image classification tasks, the loss reverts to a monotonic regime later in training. We hypothesize that this is caused by the development of in-context learning sufficient to improve n-gram sequence predictions. We observe correlational evidence in the $n$-gram sequence loss over increasing token indices and training steps in Pythia 12B (Fig. 7), where in-context learning seems to emerge in the same training step where the non-monotonic regime ends. Fascinatingly, smaller models seem to resume the standard ‘U’-shaped loss pattern in the later portions of training. [^16] Arguably, this pattern applies to models of all sizes on unigram sequences, but the tiny increases in loss for the larger models are within the margin of error for these experiments.
@@
We do not observe non-monoticity in the higher order $n$-gram sequence loss curves. However, the 3-gram loss plateaus at a point consistent with the non-monoticity observed in 1- and 2-grams, suggesting that the effect could be present to a lesser extent.
@@ 5 Conclusion
We propose two criteria that operationalize what it means for models to exploit moments of a given order, then describe methods of generating synthetic data that test whether a network satisfies both criteria, using theoretically grounded approaches relying on optimal transport theory and the principle of maximum entropy. We extend our analysis to discrete sequences by proving an equivalence between $n$-gram statistics and statistical moments.
We find new compelling empirical evidence that neural networks learn to exploit the moments of their input distributions in increasing order, and further find “double descent” in the degree to which LMs are able to model sequences sampled from low-order data statistics, driven by in-context learning on longer sequences. Our contributions strengthen the case for the distributional simplicity bias (DSB), refine our understanding of how DSB influences early learning dynamics, and provide a foundation for further investigations into DSB.
Acknowledgements
@@ Appendix B CIFARNet dataset
We ensured class balance by randomly sampling 20K images from each synset. Images were directly resized to $64\times 64$ resolution without center cropping.
@@ C.5 Street View Housing Numbers
Figure 11: The Street View Housing Numbers dataset is somewhat of an outlier in that none of the models ever exceed random baseline accuracy on 1st order synthetic images. We hypothesize this is because of the extreme diversity of colors, fonts, and background textures in SVHN, which make “simple” first order features less discriminative for classifying digits. We also found it necessary to use a smaller learning rate to achieve convergence on this dataset (Footnote 14).
@@ C.6 Pythia Language Models
Figure 12: (top) Average cross-entropy loss of Pythia models evaluated on 1- through 4-gram sequences where the 3- and 4-gram models are evaluated on a subset of the Pile, (bottom) KL divergence between the predictions of our $n$-gram language models and the predictions of Pythia checkpoints ($N=4096$.)
@@
Figure 13: Effects of fast and slow learning rate warmup on n-gram sequence loss and KL divergence, Pythia 14M and 70M ($N=4096$.)
@@
Figure 14: N-gram sequence loss over Pythia sizes and random seeds (4096 sequences sampled at each step, 9 seeds for Pythia 14M, 70M and 160M, 4 seeds for Pythia 410M.)
Figure 15: N-gram KL divergence over Pythia sizes and random seeds (4096 sequences sampled at each step, 9 seeds for Pythia 14M, 70M and 160M, 4 seeds for Pythia 410M.)
Appendix D First order optimal transport under a boundary constraint
Algorithm 1 Optimal constrained mean shift
1:Input vector $\mathbf{x}\in[0,1]^{n}$
2:Desired mean $m\in[0,1]$
3:Sort the coordinates of $\mathbf{x}$
4:$\bar{x}\leftarrow\sum_{i=1}^{n}x_{i}$
5:$\mathbf{y}\leftarrow\mathbf{0}_{n}$
6:for $i\in 1\ldots n$ do
7:  $y_{i}\leftarrow x_{i}+m-\bar{x}$
8:  if $y_{i}>1$ then
9:   $\bar{x}\leftarrow\sum_{j=i}^{n}x_{j}$
10:   $y_{i}\leftarrow 1$
11:   $m\leftarrow\frac{nm-i}{n-i}$
12:Put coordinates of $\mathbf{y}$ in their original order
13:return $\mathbf{y}$
We would like to surgically change the mean of a set of images while keeping their pixel intensities constrained to the range $[0,1]$. The least-squares optimal algorithm for this task is described in Alg. 1, and we prove its correctness in the following theorem.
Theorem D.1.
Let $\mathbf{x}$ be a vector in $[0,1]^{n}$ and let $m\in[0,1]$ be a desired mean. Then the optimization problem
$\min_{\mathbf{y}\in[0,1]^{n}}\|\mathbf{x}-\mathbf{y}\|_{2}^{2}\quad\mathrm{s.t.}\>\frac{1}{n}\sum_{i=1}^{n}y_{i}=m$
has a unique solution given by Algorithm 1.
Proof.
Let $\bar{x}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$. If $\bar{x}=m$, we immediately have the optimal solution $\mathbf{y}^{*}=\mathbf{x}$, because our constraints are already satisfied and we achieve zero loss by leaving $\mathbf{x}$ unchanged.
We can reduce the case where $m<\bar{x}$ to the case where $\bar{x}<m$ (or vice versa) by exploiting the reflection-symmetry of the problem. Specifically, if the solution to the analogous problem in $\mathbf{x}^{\prime}$ and $m^{\prime}$, where $\mathbf{x}^{\prime}=\frac{1}{2}-\mathbf{x}$ and $m^{\prime}=\frac{1}{2}-m$, is $\mathbf{y}^{\prime}$, then the solution to the original problem is $\mathbf{y}^{*}=\frac{1}{2}-\mathbf{y}^{\prime}$. This is due to the reflection-invariance of the Euclidean distance, the linearity of our mean constraint, and the fact that reflecting about $\frac{1}{2}$ cannot move points in $[0,1]^{n}$ outside $[0,1]^{n}$. Because of this symmetry, in what follows we will assume without loss of generality that $\bar{x}<m$.
Note also the optimal solution must have the property that $\forall i:x_{i}\leq y_{i}$. Assume for the sake of contradiction that $x_{i}>y_{i}$ for some $i$. Then we can improve upon this solution by increasing $y_{i}$ by some $\epsilon>0$, and compensating for this by moving another entry $y_{j}$ for which $x_{j}<y_{j}$ closer to its original value by $\epsilon$.
Setting up the Lagrangian.
Using the Karush-Kuhn-Tucker conditions, we encode the problem with the Lagrangian
$\mathcal{L}(\mathbf{y},\lambda,\boldsymbol{\mu},\boldsymbol{\nu})=\sum_{i=1}^{n}(y_{i}-x_{i})^{2}+\underbrace{\lambda\Big{(}\frac{1}{n}\sum_{i=1}^{n}y_{i}-m\Big{)}}_{\text{mean constraint}}-\underbrace{\sum_{i=1}^{n}\mu_{i}y_{i}+\sum_{i=1}^{n}\nu_{i}(y_{i}-1).}_{\text{inequality constraints}}$ (4)
Differentiating $\mathcal{L}$ with respect to $y_{i}$ yields the stationarity condition
$\frac{\partial\mathcal{L}}{\partial y_{i}}=2(y_{i}-x_{i})+\frac{\lambda}{n}-\mu_{i}+\nu_{i}=0.$ (5)
The KKT complementary slackness condition requires that $\mu_{i}y_{i}=0$ and $\nu_{i}(y_{i}-1)=0$ for each $i$. This implies that $\mu_{i}$ must be zero if $y_{i}>0$, and $\nu_{i}$ must be zero if $y_{i}<1$. For each $i$ where $y_{i}<1$, we can use Eq. 5 and complementary slackness to write $y_{i}$ as $x_{i}-\frac{\lambda}{2n}$.
Putting it all together.
Assume $\mathbf{x}$ and $\mathbf{y}$ are written in a basis that ensures the coordinates of $\mathbf{x}$ are sorted in descending order, so that $x_{1}\geq x_{2}\geq\ldots\geq x_{n}$. Our problem is invariant to permutation of indices, so this does not affect the solution.
We can now solve for $y_{1}$, the final position of the largest coordinate, in the following way. Suppose that $y_{1}<1$. Then we have $\forall i:y_{i}<1$, and the mean constraint can be written as $\frac{1}{n}(\sum_{i=1}^{n}x_{i})-\frac{\lambda}{2n}=m$. This allows us to solve for all $y_{i}$:
$y_{i}=x_{i}-\frac{\lambda}{2n}=x_{i}+m-\frac{1}{n}(\sum_{i=1}^{n}x_{i}).$ (6)
Note Eq. 6 may “overshoot” and violate the inequality constraint $y_{i}\leq 1$. If it does, then we know our supposition is false and $y_{1}=1$. If it does not violate the constraint, then it must be optimal because it is also the solution to the relaxed version of this problem without the $[0,1]$ constraint. In the latter case, we are done.
Given that $y_{1}=1$, the subproblem of solving for $y_{2},\ldots,y_{n}$ is a smaller instance of the original problem: the target mean for these coordinates is $m^{\prime}=\frac{nm-1}{n-1}$. We can recursively apply this reasoning to solve for all other $y_{i}$. This procedure coincides with Algorithm 1. ∎
Appendix E Derivation of the Conrad distribution
Theorem E.1.
Among all distributions supported on $[0,1]$ with desired mean $m\neq\frac{1}{2}$, the Conrad distribution with density $p(x)=\frac{b\exp(-bx+b)}{\exp(b)-1}$ has maximum entropy, where the parameter $b\neq 0$ is chosen to satisfy the equation $m=-\frac{b-\exp(b)+1}{b(\exp(b)-1)}$. In the special case of $m=\frac{1}{2}$, the maximum entropy distribution is $\mathrm{Unif}(0,1)$.
Proof.
[…]
$H(q)$ $\leq H(q,p)$ (inequality of entropy and cross-entropy)
$=\mathbb{E}_{q}[\log p(x)]$ (definition of cross-entropy)
$=-a-bm$ (definition of $p(x)$ and linearity)
$=H(p)$ (QED)
We can now analytically solve for $a$ in terms of $b$. Integrating $p(x)$ from 00 to $1$ yields $-\frac{\exp(-a-b)}{b}+\frac{\exp(-a)}{b}=1$. Solving for $a$ we get $a=-b+\log\big{(}\frac{\exp(b)-1}{b}\big{)}$, which when plugged back into the original formula gives us $p(x)=\frac{b\exp(-bx+b)}{\exp(b)-1}$.
[…]
$\lim_{b\rightarrow 0}-\frac{b-\exp(b)+1}{b(\exp(b)-1)}=\lim_{b\rightarrow 0}\frac{\exp(b)-1}{b\exp(b)+\exp(b)-1}=\lim_{b\rightarrow 0}\frac{\exp(b)}{b\exp(b)+2\exp(b)}=\frac{1}{2}.$ (7)
The maximum entropy distribution supported on $[0,1]$ with no mean constraint is known to be $\mathrm{Unif}(0,1)$. Since it happens to have the mean $\frac{1}{2}$, we may conclude that the Conrad distribution approaches $\mathrm{Unif}(0,1)$ as $b$ approaches 00. ∎
Appendix F Discrete domain proofs